OpenStudy (tiffany_rhodes):

Find the solution of the given initial value problem: y' - y= 6e^2t. y(0) = 1.

OpenStudy (tiffany_rhodes):

I figured you would use the product rule in reverse

OpenStudy (anonymous):

\[y'-y=6e^{2t}\] Find your integrating factor first: \[\mu(t)=\exp\left(\int (-1)~dt\right)=e^{-t}\] Multiply both sides of the eq. by \(\mu\): \[e^{-t}y'-e^{-t}y=6e^{t}\] As you mentioned, the left side is the result of differentiation via product rule: \[\frac{d}{dt}\left[e^{-t}y\right]=6e^{t}\] Anti-differentiate both sides: \[e^{-t}y=6\int e^{t}~dt\] And so on. You'll have a constant of integration on the right side; plug in the initial conditions to solve for it.

OpenStudy (tiffany_rhodes):

thanks @SithsAndGiggles !

OpenStudy (anonymous):

yw